### v1

@@ 5 Sensibility Analysis
PathMNIST Mean 91.3 17.9($\downarrow$73.4) 16.4($\downarrow$74.9) 17.1($\downarrow$74.2) 15.6(normal-$\downarrow$75.7)
Median 90.8 22.4($\downarrow$68.4) 7.2(normal-$\downarrow$83.6) 12.7($\downarrow$78.1) 19.3($\downarrow$71.5)
Gaussian 92.1 35.5($\downarrow$56.6) 11.0(normal-$\downarrow$81.1) 18.0($\downarrow$74.1) 17.7($\downarrow$74.4)
DermaMNIST Mean 73.8 45.8($\downarrow$28.0) 31.9($\downarrow$41.9) 23.9($\downarrow$49.9) 21.5(normal-$\downarrow$52.3)
Median 75.1 62.6($\downarrow$12.5) 34.6($\downarrow$40.5) 19.9($\downarrow$55.2) 17.8(normal-$\downarrow$57.3)
Gaussian 74.7 40.1($\downarrow$34.6) 33.1($\downarrow$41.6) 25.8(normal-$\downarrow$48.9) 40.6($\downarrow$34.1)
OctMNIST Mean 77.7 25.7($\downarrow$52.0) 28.3($\downarrow$49.4) 23.8(normal-$\downarrow$53.9) 29.6($\downarrow$48.1)
Median 79.6 23.5($\downarrow$56.1) 25.7($\downarrow$53.9) 23.6($\downarrow$56.0) 22.6(normal-$\downarrow$57.0)
Gaussian 77.7 22.9(normal-$\downarrow$54.8) 26.1($\downarrow$51.6) 23.1($\downarrow$54.6) 29.2($\downarrow$48.5)
[…]
Median 90.9 38.2($\downarrow$52.7) 62.7($\downarrow$28.2) 50.9($\downarrow$40.0) 28.4(normal-$\downarrow$62.5)
Gaussian 91.1 45.1($\downarrow$46.0) 70.8($\downarrow$20.3) 51.4($\downarrow$39.7) 29.8(normal-$\downarrow$61.3)
@@
Resistance to Cropping. As we consistently highlighted, medical data exhibits greater sparsity compared to conventional data. For instance, when processing microscopic images, physicians or researchers typically reduce the proportion of the background through cropping and magnifying details to facilitate analysis. In this section, we further validate the effectiveness of our SALM after cropping. We calculate the difference in test accuracy on the protected data before and after cropping. To minimize the impact of cropping itself on performance, we add the difference in test accuracy on the clean data before and after cropping, which we denote as the Gap value. The results in Table 5 show after cropping, the performance gap of TAP and EM significantly narrows. This finding confirms that color space disturbances by these methods are random. If perturbations are concentrated in non-feature areas like transparent slides, cropping might result in a loss of protection. Conversely, SALM focuses on pixels that significantly contribute to the model, which ensures that removing the background does not impact the regions with more protective noise and thus preserves the protection.
6 Utility Analysis
@@
Table 5: Image similarity scores of different methods.
@@
SALM 90.9 82.5 14.9 5.3 1.2normal-$\uparrow$
@@
Table 5: Image similarity scores of different methods.
Figure 3: The visualization of the origin image and protected image with noise and their corresponding segmentation results.
Medical imaging plays an important role in the diagnosis of a wide range of diseases. Imaging technologies enable doctors to view more detailed structures, thus facilitating more accurate disease diagnoses. Consequently, it is imperative to not only thwart unauthorized training effectively but also to ensure that perturbations in images do not impair the judgment of doctors. Previous studies indicate that when the perturbed radius is smaller than 8, it doesn’t affect visual judgment. However, the implications in the real world of medical data call for an in-depth assessment of image usability, which is the focus of this section.
We start our analysis by evaluating image similarity between the original image and the one protected. Specifically, we utilized the Structural Similarity Index Measure (SSIM), which assesses luminance, contrast, and structure, along with various Hash methods focusing on low-frequency information. The results in Tab 5 reveal that SYN has the highest scores. This is likely due to its structure, consisting of square regions of differing colors with minimal distinctions within these areas. However, compared with other methods, despite being based on feature perturbations, our method does not substantially affect images. This underscores our approach’s ability to effectively preserve image utility.
Furthermore, physicians often use pre-trained models to segment blood vessels or tissues for auxiliary. Therefore, it’s essential that the protection applied to images does not compromise segmentation accuracy. The results in Fig. 3 show that the segmentation results of the protected image are unaffected. To further validate this, we compared the segmentation masks and observed that in 224x224 pixel images, only 0.4% of the pixels showed any difference from the originals. These pixels had inherently low confidence in the original image segmentation, leading to minor perturbations and changes in their results. Overall, the impact of SALM on segmentation performance is negligible, with the protected images still able to guide toward accurate auxiliary results.
7 Case Study: Large-scale Cooperative Dataset
Constructing a large-scale dataset in the real world often requires collaboration among many researchers or institutions. Before the final integration for release, the publisher has no obligation to protect the integrity of the overall dataset. Consequently, it’s crucial for each collaborating entity to implement robust methods to protect their respective data contributions. Previous experiments have assumed access to all data, which is unlikely during the dataset construction phase. Therefore, this section will investigate the efficacy of our approach when only partial data is available, and there is complete unawareness of the remaining data. We set up two scenarios for validation: class-wise combine and sample-wise combine.
@@ 7 Case Study: Large-scale Cooperative Dataset
7.1 Class-wise Combined Dataset
We initially conducted a study on the scenario of the class-wise combination. To better explain the experimental setting, we introduce the concept of “support classes,” which is utilized to augment the existing single-class data. We selected two kinds of support classes: medical classes and general classes. The framework is shown in Fig. 4. First, we chose melanoma from the DermaMNIST as a protected target and selected all classes within the PathMNIST as support classes for our experiments. The melanoma data was combined with PathMNIST to create a ten-class dataset while only preserving the SALM noise generated for the melanoma. Subsequently, to broaden the selection range of support classes, we also selected some general image classes (e.g., airplanes) as support classes.
Following the generation of class-specific noise, we utilized the same experimental settings as in previous tests. Using PathMNIST for support classes, the generated noise resulted in a clean test accuracy of 9.9%. Noise generated with general image classes from ImageNet [12] as support classes led to an accuracy of 8.1%. The choice of either type of support class does not affect the noise performance. This demonstrates the flexible applicability of our method.
7.2 Sample-wise Combined Dataset
@@
The results in Table 6 show a rapid decrease in effectiveness when less than 100% of the data is selected for SALM application, whatever the $K$ is. Surprisingly, applying SALM to as much as 80% of the data results in a negligible effect. Previous studies have shown that both error-minimizing noise and error-maximizing noise have similar limitations in DNNs [25, 56].
@@
To further illustrate this phenomenon, we take 80% as an example and plot the learning curves of ResNet-18 in the following scenarios: 1) Models trained on 20% clean data; 2) Models trained on 80% data processed by SALM; 3) Models trained on a mixture of these two types of data. The results are shown in Figure 5 left. It is observed that data processed by SALM at 80% remains unlearnable for the model, yet the mere 20% of clean data enables the model to achieve excellent performance. The superior performance trained on mixed data arises solely from the clean data contained within. Furthermore, the visualization results in Fig. 5 reveal that the clean data remains mixed in 2D space, whereas the data processed by SALM is almost linearly separable, indicating its unlearnability. Therefore, in the case of sample-wise combined datasets, our method also demonstrates outstanding performance and robust scalability. Overall, our method provides a way for every collaborating entity within large datasets to protect their interests and ensure each data donor’s contribution will not be misused.
8 Conclusion
Appendix 0.A Datasets Details
Table 7: The details about the datasets we utilized for our experiments. The default size is 28$\times$28, and the corresponding 224 version is only different in size.
Dataset Modality Task Scale
PathMNIST Colon Pathology Multi-Class(9) 107180
ChestMNIST Chest X-Ray Multi-Label (14)&Binary-Class (2) 112,120
DermaMNIST Dermatoscope Multi-Class (7) 10,015
OCTMNIST Retinal OCT Multi-Class (4) 109,309
PneumoniaMNIST Chest X-Ray Binary-Class (2) 5,856
RetinaMNIST Fundus Camera Ordinal Regression (5) 1,600
BreastMNIST Breast Ultrasound Binary-Class (2) 780
BloodMNIST Blood Cell Microscope Multi-Class (8) 17,092
TissueMNIST Kidney Cortex Microscope Multi-Class (8) 236,386
OrganAMNIST Abdominal CT Multi-Class (11) 58,830
OrganCMNIST Abdominal CT Multi-Class (11) 23,583
OrganSMNIST Abdominal CT Multi-Class (11) 25,211
Appendix 0.B More Implementation Details
0.B.1 Hardware Details
The whole experiments are conducted on an Ubuntu 20.04.6 LTS. Experiments For all datasets are conducted on 1 GPU (NVIDIA RTX 4090 24GB) and 16 CPU cores (Intel(R) Xeon(R) Platinum 8352V CPU @ 2.10GHz). Python 3.10.8 and Pytorch 2.1.2 with CUDA 12.1 are used for all the implementations.
0.B.2 Experiments details
PGD. For the PGD process, we set PGD as a random start with total steps $K_{a}=10$, step size $\alpha_{u}=1.6$, and radius $\rho_{u}=8$. Such parameters not only balance protection time and protection performance but also ensure the utility of the data.
Testing. We employ four models introduced in Sec 4.1 for our experiments. All models were trained for 20,000 iterations, and then the accuracy results are reported on their respective data types. For datasets with a resolution of 28x28, the batch size was set to 128, while for datasets of 224x224 resolution, the batch size was reduced to 64. Additionally, for VGG-11, images were padded to a size of 32x32. All other specific parameters were consistent with those used during the noise generation phase.
Appendix 0.C Baseline Details
0.C.1 Target Adversarial Poisoning
This method uncovers the further potential of adversarial examples, originally intended for attacking pre-trained models. Adversarial examples also demonstrate the powerful effects in data poisoning. Specifically, by utilizing the empirical loss maximization problem in place of the bi-level problem according to Eq (7), the poison generation problem is transformed into an adversarial example issue.
$\max_{\delta\in\mathcal{S}}\left[\sum_{\left(x_{i},y_{i}\right)\in\mathcal{T}}\mathcal{L}\left(F\left(x_{i}+\delta_{i};\theta^{*}\right),y_{i}\right)\right]$ (7)
0.C.2 Error Minimizing Noise
We have already introduced EM in the main text. More specifically, we utilize RN-18 as the source model for generation according to Eq. (2). But here we employed the iterative generation framework within REM to achieve enhanced protection effects. The PGD method is employed for solving the inner minimization problem, and the settings are shown in Appendix 0.B.2.
0.C.3 Synthetic Perturbation
This approach induces data to become linearly separable by incorporating multiple square high-frequency patches into images. Specifically, the method initially generates some vertex coordinates in the form of hypercubes in high-dimensional space, then repeats each dimension to introduce local correlations, and finally scales down. For the perturbation radius, we similarly set it to 8/255.
Appendix 0.D More Results
We present additional clean test accuracy results for the 224x224 resolution in Tab 8. The symbol $\downarrow$ in the following context indicates a decrease in accuracy compared to the clean test accuracy.
Dataset Clean AdvT EM SP SALM($K=10$)
[…]
Figure 6: The prediction matrix of models trained on data with three processed classes: CAS, CAE and LYM.
Figure 6: The prediction matrix of models trained on data with three processed classes: CAS, CAE and LYM.
Figure 7: Clean test accuracy(%) of DNNs trained on three selected clean and poisoned classes.
0.D.1 Partial-class SALM.
Here, we assume that a complete dataset is already available and can be fully accessed and processed, as an ideal scenario. This serves as the theoretical foundation for our further case studies in real-world situations. In medical datasets, specific classes frequently attract more attention than others. We selected three classes in PathMNIST: CAS, Colorectal Adenocarcinoma Epithelium (CAE), and Lymphocytes (LYM). SALM was applied to all samples within these selected classes. The results in Fig. 7 illustrate that the selected classes become unlearnable. Crucially, as shown in Fig. 7, the misclassification occurred only in the selected classes. These findings underscore the adaptability of our method. Assuming sufficient computational resources and unrestricted access, in a collaborative dataset, researchers can secure the classes they gather. Even in the final combined dataset, these segments retain limited learning utility for their specific categories, thus protecting the interests of contributors.
0.D.2 Case Study: Class-wise Combined Dataset.
In our case study, we solely conducted experiments on the melanoma category. We replicated this procedure for all categories within DermaMNIST and then aggregated the results. For the selection of support classes, we chose PathMNIST, BloodMNIST, and CIFAR-10, randomly selecting five support classes. On the dataset contaminated and subsequently merged in this manner, the test accuracy of the trained RN-18 was merely 10.3%, which is almost equivalent to the protective effect observed when accessing the entire dataset.
@@ Appendix 0.D More Results
Figure 8: More visualization results of the origin image and the corresponding noise and protected image.
Figure 9: More visualization results of the origin image and protected image with noise and their corresponding segmentation results.

### v2

@@ 4 Experiments and Results
PathMNIST Mean 91.3 17.9($\downarrow$73.4) 16.4($\downarrow$74.9) 17.1($\downarrow$74.2) 15.6($\downarrow$75.7)
Median 90.8 22.4($\downarrow$68.4) 7.2($\downarrow$83.6) 12.7($\downarrow$78.1) 19.3($\downarrow$71.5)
Gaussian 92.1 35.5($\downarrow$56.6) 11.0($\downarrow$81.1) 18.0($\downarrow$74.1) 17.7($\downarrow$74.4)
DermaMNIST Mean 73.8 45.8($\downarrow$28.0) 31.9($\downarrow$41.9) 23.9($\downarrow$49.9) 21.5($\downarrow$52.3)
Median 75.1 62.6($\downarrow$12.5) 34.6($\downarrow$40.5) 19.9($\downarrow$55.2) 17.8($\downarrow$57.3)
Gaussian 74.7 40.1($\downarrow$34.6) 33.1($\downarrow$41.6) 25.8($\downarrow$48.9) 40.6($\downarrow$34.1)
OctMNIST Mean 77.7 25.7($\downarrow$52.0) 28.3($\downarrow$49.4) 23.8($\downarrow$53.9) 29.6($\downarrow$48.1)
Median 79.6 23.5($\downarrow$56.1) 25.7($\downarrow$53.9) 23.6($\downarrow$56.0) 22.6($\downarrow$57.0)
Gaussian 77.7 22.9($\downarrow$54.8) 26.1($\downarrow$51.6) 23.1($\downarrow$54.6) 29.2($\downarrow$48.5)
ChestMNIST224 Mean 91.4 37.7($\downarrow$53.7) 72.5($\downarrow$18.9) 58.2($\downarrow$33.2) 27.0($\downarrow$64.4)
Median 90.9 38.2($\downarrow$52.7) 62.7($\downarrow$28.2) 50.9($\downarrow$40.0) 28.4($\downarrow$62.5)
Gaussian 91.1 45.1($\downarrow$46.0) 70.8($\downarrow$20.3) 51.4($\downarrow$39.7) 29.8($\downarrow$61.3)
@@
5 Conclusion
Appendix A More Related Work
Medical Data Protection. As information technology advances, digital technologies are increasingly integrated into medicine, affecting both clinical treatment and scientific research (Senbekov et al., 2020; Rajpurkar et al., 2017; Johnson et al., 2019; Irvin et al., 2019; Pedrosa et al., 2019; Bilic et al., 2023; Rasmy et al., 2021; Zhang et al., 2022; Kelly et al., 2019). However, these developments also present significant risks and challenges regarding patient privacy, medical information breaches occur frequently around the world (Alpert, 2003; Kayaalp, 2018; Price & Cohen, 2019; Murdoch, 2021). Inadequate data protection can lead to substantial harm through the leakage of personal information, such as the disclosure of sensitive health conditions like HIV, potentially leading to social isolation and psychological disorders (Gostin, 2000). Furthermore, incidents of data breaches may also reduce patients’ trust in medical research institutions, making them reluctant to share their data (Koh et al., 2011; Trinidad et al., 2020). Earlier studies focused on robust physical protection measures, like using encrypted storage devices (Heurix & Neubauer, 2011; Akinyele et al., 2011; Sun et al., 2020), establishing firewalls (Barrows Jr & Clayton, 1996; Liu et al., 2012), and secure communication transmission modes (Gong et al., 2015; Li et al., 2016), to safeguard data. Liu & Li (2018) introduced a clustering method based on $K$-anonymity algorithm as the building block of privacy-preserving for medical devices’ data. However, as collaboration among various research institutions intensifies and open-source data sharing on the internet becomes more crucial, these methods become less applicable. Baowaly et al. (2019); Yoon et al. (2020) attempted to use generated data to reduce granularity and thus protect privacy, but this method involves a significant trade-off between information loss and protection efficacy. However, our method remains robust under various conditions during the construction process of real-world datasets and effectively balances high data usability with consistent protection.
Appendix B SALM Algorithm
Algorithm 1 Training SALM generator and generating noise.
  Input: Training data set $\mathcal{T}$, Training steps $M$, PGD parameters $\alpha_{u}$ and $\rho_{u}$, transformation distribution $T$, the sampling number $J$ for gradient approximation.
  Initialization: source model parameter $\theta$, $\delta^{u}$.
  // Following Eq. (3)
  for $i$ in $1,\cdots,M$ do
     Sample minibatch $(x,y)\sim\mathcal{T}$, sample transformation $t_{j}\sim T$
     Calculate $g_{k}\leftarrow\frac{1}{J}\sum_{j=1}^{J}\frac{\partial}{\partial\delta^{u}}\ell(f^{\prime}_{\theta}(t_{j}(x+\delta^{u})),y)$
     Determine $k$-th percentile value $g^{0}(k)$ in $g_{k}$
     for each element $g_{ij}$ in $g_{k}$ do
        if $g_{ij}\geq g^{0}(k)$ then
           Set $M_{ij}=1$
        else
           Set $M_{ij}=0$
        end if
     end for
     Apply local mask: $\mathcal{P}(g_{k},k)=M\odot g_{k}$  // Following Eq. (5)
     Update $\delta^{u}\leftarrow\prod_{\|\delta\|\leq\rho_{u}}\left(\delta^{u}-\alpha_{u}\cdot\mathrm{sign}(\mathcal{P}(g_{k},k))\right)$
     Update source model parameter $\theta$ based on minibatch $(t(x+\delta^{u}),y)$
  end for
  Output: SALM noise generator $f^{\prime}_{\theta}$, SALM noise $\delta^{u}$
Appendix C Experiments Setup
Datasets. Our research involved extensive experiments on MedMNIST (Yang et al., 2021), which is a comprehensive collection of standardized biomedical images spanning 12 distinct datasets. MedMNIST includes the major modalities in medical imaging. This dataset showcases diverse data scales and a range of tasks. Furthermore, to more effectively assess our method’s performance on large, real-world datasets, we also do experiments on the 224x224 datasets from MedMNIST-v2 (Yang et al., 2023).
Model and Implementation Details. We selected the following well-known models to evaluate our method against benchmark approaches: VGG-11 (Simonyan & Zisserman, 2014), ResNet-18/50 (He et al., 2016) and DenseNet-121 (Huang et al., 2017). For the SALM generator, we select RN-18 as the source model to generate noise. We set the perturbation noise $\rho_{u}=8/255$ as default. We choose an SGD optimizer for both noise generation and training. Both of their weight decay are set to 5e-4, and momentum is set to 0.9. The initial learning rate is set to 0.1 with a decay rate of 0.1.
Baselines. Since there’s no previous method specifically designed for medical data, we compare our Sparsity-Aware Local Masking (SALM) with the existing SoTA methods in the general image domain. The baseline methods are as following: Target Adversarial Poisoning (TAP) (Fowl et al., 2021), Error-Minimizing noise (EM) (Huang et al., 2021), Synthesized Perturbation (SP) (Yu et al., 2022). None of these three methods impose restrictions on the range of pixels to be perturbed. TAP and EM both target the training loss and craft noise using the gradient information to trigger maximum error or trick models to overfit. SC, on the other hand, imposes hand-crafted linear-separable noise into the data, thereby leading the model to learn only simple noise-label correlation.
Appendix D More Experiments
Dataset Clean AdvT EM SP SALM($K=10$)
DermaMNIST224 74.8 19.2($\downarrow$55.6) 11.4($\downarrow$63.4) 10.0($\downarrow$64.8) 9.3($\downarrow$65.5)
OCTMNIST224 76.2 32.3($\downarrow$43.9) 14.6($\downarrow$61.6) 20.6($\downarrow$55.6) 25.8($\downarrow$50.4)
PneumoniaMNIST224 84.1 66.7($\downarrow$17.4) 67.2($\downarrow$16.9) 61.6($\downarrow$22.5) 57.4($\downarrow$26.7)
RetinaMNIST224 51.0 27.6($\downarrow$23.4) 21.1($\downarrow$29.9) 10.5($\downarrow$40.5) 9.1($\downarrow$41.9)
BreastMNIST224 80.1 35.1($\downarrow$45.0) 40.0($\downarrow$40.1) 40.2($\downarrow$39.9) 37.8($\downarrow$42.3)
BloodMNIST224 96.0 16.3($\downarrow$80.3) 16.6($\downarrow$79.4) 7.0($\downarrow$89.0) 6.3($\downarrow$89.7)
TissueMNIST224 68.4 18.5($\downarrow$49.9) 23.6($\downarrow$44.8) 16.7($\downarrow$51.7) 19.1($\downarrow$49.3)
OrganAMNIST224 95.2 85.8($\downarrow$9.4) 83.5($\downarrow$11.7) 85.3($\downarrow$9.9) 82.1($\downarrow$13.1)
OrganCMNIST224 91.6 82.3($\downarrow$9.3) 69.7($\downarrow$21.9) 72.2($\downarrow$19.4) 69.3($\downarrow$22.3)
OrganSMNIST224 77.8 47.8($\downarrow$30.0) 44.6($\downarrow$33.2) 65.2($\downarrow$12.6) 42.1($\downarrow$35.7)
@@ Appendix D More Experiments
SALM 90.9 82.5 14.9 5.3 1.2$\uparrow$
Table 6: Image similarity scores of different methods.
@@
D.1 Resistance to Cropping.
As we consistently highlighted, medical data exhibits greater sparsity compared to conventional data. For instance, when processing microscopic images, physicians or researchers typically reduce the proportion of the background through cropping and magnifying details to facilitate analysis. In this section, we further validate the effectiveness of our SALM after cropping. We calculate the difference in test accuracy on the protected data before and after cropping. To minimize the impact of cropping itself on performance, we add the difference in test accuracy on the clean data before and after cropping, which we denote as the Gap value. The results in Table 6 show after cropping, the performance gap of TAP and EM significantly narrows. This finding confirms that color space disturbances by these methods are random. If perturbations are concentrated in non-feature areas like transparent slides, cropping might result in a loss of protection. Conversely, SALM focuses on pixels that significantly contribute to the model, which ensures that removing the background does not impact the regions with more protective noise and thus preserves the protection.
D.2 Similarity Compare.
Although previous studies (Liu et al., 2023d; Huang et al., 2021; Fu et al., 2022; Fowl et al., 2021; Yu et al., 2022) have demonstrated that perturbation radii as small as 8 or even 16 do not affect human visual perception, considering the sensitivity of medical images, we still assessed the similarity. Specifically, we utilized the Structural Similarity Index Measure (SSIM), which assesses luminance, contrast, and structure, along with various Hash methods focusing on low-frequency information. The results in Table 6 reveal that SYN has the highest scores. This is likely due to its structure, consisting of square regions of differing colors with minimal distinctions within these areas. However, compared with other methods, despite being based on feature perturbations, our method does not substantially affect images. This underscores our approach’s ability to effectively preserve image utility.
@@
Appendix E Case Study
E.1 Case 1: Class-wise Combined Dataset
We initially conducted a study on the scenario of the class-wise combination. To better explain the experimental setting, we introduce the concept of “support classes,” which is utilized to augment the existing single-class data. We selected two kinds of support classes: medical classes and general classes. The framework is shown in Figure 5. First, we chose melanoma from the DermaMNIST as a protected target and selected all classes within the PathMNIST as support classes for our experiments. The melanoma data was combined with PathMNIST to create a ten-class dataset while only preserving the SALM noise generated for the melanoma. Subsequently, to broaden the selection range of support classes, we also selected some general image classes (e.g., airplanes) as support classes.
Following the generation of class-specific noise, we utilized the same experimental settings as in previous tests. Using PathMNIST for support classes, the generated noise resulted in a clean test accuracy of 9.9%. Noise generated with general image classes from ImageNet (Deng et al., 2009) as support classes led to an accuracy of 8.1%. The choice of either type of support class does not affect the noise performance. This demonstrates the flexible applicability of our method.
E.2 Case 2: Sample-wise Combined Dataset
@@
The results in Table 7 show a rapid decrease in effectiveness when less than 100% of the data is selected for SALM application, whatever the $K$ is. Surprisingly, applying SALM to as much as 80% of the data results in a negligible effect. Previous studies have shown that both error-minimizing noise and error-maximizing noise have similar limitations in DNNs (Huang et al., 2021; Shan et al., 2020).
@@
To further illustrate this phenomenon, we take 80% as an example and plot the learning curves of ResNet-18 in the following scenarios: 1) Models trained on 20% clean data; 2) Models trained on 80% data processed by SALM; 3) Models trained on a mixture of these two types of data. The results are shown in Figure 6 left. It is observed that data processed by SALM at 80% remains unlearnable for the model, yet the mere 20% of clean data enables the model to achieve excellent performance. The superior performance trained on mixed data arises solely from the clean data contained within. Furthermore, the visualization results in Figure 6 reveal that the clean data remains mixed in 2D space, whereas the data processed by SALM is almost linearly separable, indicating its unlearnability. Therefore, in the case of sample-wise combined datasets, our method also demonstrates outstanding performance and robust scalability. Overall, our method provides a way for every collaborating entity within large datasets to protect their interests and ensure each data donor’s contribution will not be misused.
@@
Figure 7: Some visualization results of the origin image and the corresponding noise and protected image.